A firm has four empty seats to fill on its board. The total number of shares outstanding is 473,000 shares and Annan owns 90,000 shares. Can Annan guarantee that he will get on the board, if cumulative voting method is used? A. No, he doesn't have enough shares to guarantee membership on the board. B. No, he needs a minimum of 99,601 shares to guarantee membership on the board. C. Yes, he owns more than enough number of shares. D. Yes, he needs only 87,201 shares and he has more than that.
Answer #1

Total number of shares of common stock outstanding = 473,000 shares

For election to four seats on the board of directors, if the firm uses cumulative voting to elect its board,

Total available votes = 473,000 *4 = 1,892,000

The four candidates with the highest number of votes will be elected to the board, therefore the minimum number of votes needed to ensure election of one member to the board = [Total available votes / (number of seats on the board of directors+1)] +1   (one is added to avoid any tie)

= [1,892,000 / (4+1)] +1

= 378,400 + 1

= 378,401

Now remaining number of votes are = 1,892,000 - 378,401 = 1,513,599 if it is divided between four other candidates, one will definitely get less than 378,401 votes and candidate with minimum 378,401 votes will definitely win.

As Annan owns 90,000 shares, therefore he doesn't have enough shares to guarantee membership on the board

Therefore correct answer is option A. No, he doesn't have enough shares to guarantee membership on the board.